# A simple model of globally magnetized accretion discs

### v1

@@ Abstract
We present an analytic, quasi-local dynamo model for accretion discs threaded by net, vertical magnetic flux. In a simple slab geometry and ignoring stochastic mean-field dynamo effects, we calculate the large-scale field resulting from the balance between kinematic field amplification and turbulent resistive diffusion. The ability of the disc to accumulate magnetic flux is sensitive to a single parameter dependent on the ratio of the vertical resistive diffusion time to the Alfvén crossing time, and we show how the saturation levels of magnetorotational and other instabilities can govern disc structure and evolution. Under wide-ranging conditions, inflow is governed by large-scale magnetic stresses rather than internal viscous stress. We present models of such “magnetically boosted” discs and show that they lack a radiation pressure-dominated zone. Our model can account for “magnetically elevated” discs as well as instances of midplane outflow and field reversals with height that have been seen in some global simulations. Using the time-dependent features of our model, we find that the incorporation of dynamo effects into disc structure can lead to steady or episodic “magnetically arrested discs” (MADs) that maximize the concentration of magnetic flux in their central regions.
@@
pubyear: 2024pagerange: A simple model of globally magnetized accretion discsA simple model of globally magnetized accretion discs
@@ 1 Introduction
Magnetic boosting of the inflow speed as well as magnetic elevation can be critical for understanding the dynamics, spectra and stability of accretion discs in a variety of astrophysical systems. Even before these computational trends had been established, the existence of magnetically supported discs had been proposed for phenomenological reasons (Begelman & Pringle, 2007; Gaburov et al., 2012; Sądowski, 2016; Begelman & Silk, 2017; Dexter & Begelman, 2019). Yet while magnetic elevation seems to be a robust feature of simulated flows, there is to date no clear explanation of how large-scale, organized magnetic field structures are created, and what factors determine their properties.
@@
Even more extreme are so-called magnetically arrested discs (MADs) (Narayan et al., 2003; Igumenshchev, 2008), which have been studied intensively as a type of accretion flow that maximizes the magnetic flux threading a black hole, thus maximizing the potential to extract energy from the black hole’s spin (Tchekhovskoy et al., 2011). While the scaling between the accretion rate and trapped magnetic flux in a MAD was initially derived by equating the ram-pressure of free-fall accretion to the magnetic pressure of the trapped flux (Bisnovatyi-Kogan & Ruzmaikin, 1974, 1976; Narayan et al., 2003), arguing that the magnetic flux would “arrest" the flow, it has been since recognized MAD models permit continuous accretion and that the accumulation of magnetic flux in a MAD can extend to much larger radii than the black hole horizon, in effect creating a large-scale magnetosphere (McKinney et al., 2012) co-existent with the accretion flow.
@@
Despite the potential for magnetically elevated and arrested discs to explain a wide range of accretion phenomena, there remains deep uncertainty about how to establish and maintain them in the first place. A large body of work, beginning with van Ballegooijen (1989) and Lubow et al. (1994), suggests that under most conditions the flux distribution will evolve with time. In particular, Lubow et al. (1994) highlighted the challenge faced by thin discs in accumulating and retaining flux, finding that the disc will advect and retain flux only if the poloidal field lines at the disc surface make an angle with the vertical smaller than $\sim Hv_{r}/\eta$, where $H$ is the disc scale height, $v_{r}$ is the inflow speed inside the disc, and $\eta$ is the effective electrical resistivity in the disc interior, in most cases assumed to be due to turbulence. For standard viscous disc models this critical angle is $\sim H/r\ll 1$, whereas constraints imposed by the external electromagnetic environment of the disc — whether by matching the interior field to a magnotocentrifugal wind (Blandford & Payne, 1982; Königl, 1989; Li, 1995) or some other physically motivated field configuration (Lubow et al., 1994; Okuzumi et al., 2014; Takeuchi & Okuzumi, 2014) — tend to favor an angle of order unity.
@@
Various effects may help a disc to retain magnetic flux, including thickening of the disc due to magnetic pressure support, reduced resistivity in the disc surface layers (Lovelace et al., 2009; Guilet & Ogilvie, 2012, 2013), and enhanced inflow speeds due to large-scale magnetic torques (Blandford & Payne, 1982; Ferreira & Pelletier, 1993; Königl, 1989; Li, 1995). In this paper, we focus on the latter effect, arguing that enhanced inflow speeds may be a natural feature of accretion discs under many conditions, even in the standard thin-disk limit in which the disc is supported by gas pressure and the magnetorotational instability (MRI) operates as usual.
@@
To make this case — and explore its broader implications — we construct and analyze a quasi-local, time-dependent, analytic dynamo model for thin discs threaded by a net vertical magnetic field. The plan of the paper is as follows. In §2 we introduce the kinematic dynamo model (§2.1) and solve for the magnetic field structure using boundary conditions (§2.2) that connect the disc to its external environment. We use these solutions to derive disc evolution equations (§2.3), which show that magnetic stresses are at least as important for driving inflow and magnetic flux evolution as internal viscous stresses, and are often much more important. We refer to these rapidly inflowing discs as “magnetically boosted.” We conclude the section with a discussion of the reduced levels of dissipation expected inside magnetized discs (§2.4), one consequence of which is the absence of any radiation pressure-dominated inner zone (§4.1).
@@
The analysis of §2 reveals the critical importance of a parameter (labeled $q$) which essentially represents the ratio of the resistive diffusion time across the disc scale height to the Alfvén crossing time. Given $B_{z}$, this ratio mainly depends on the level of turbulence inside the disc, driven by magnetorotational and perhaps other instabilities. In §3, we discuss how this ratio might differ in weakly and strongly magnetized discs, while recognizing that the saturation levels of disc turbulence are among the least understood inputs for this or any similar model. Nevertheless, the adoption of a specific closure relation allows us to construct steady-state models of magnetically boosted accretion discs in §4, for both weakly and strongly magnetized limits. We provide a preliminary, local analysis of the stability of these models in §5.
@@
We summarize our results in §7 and expand on two further implications: the possibility of midplane outflow and field reversals within the disc, and potential relevance of our model to episodic accretion phenomena such as X-ray nova outbursts and changing-look AGN.
@@ 2 Magnetized disc model
Extensive literature exists on accretion disc dynamos, aimed at understanding both the turbulent and organized magnetic fields. Turbulent accretion disc dynamos, which have been the subject of numerous numerical and analytic studies (Kato & Yoshizawa, 1995; Vishniac & Brandenburg, 1997; Ogilvie, 2003; Gressel, 2010; Käpylä & Korpi, 2011; Gressel & Pessah, 2015; Squire & Bhattacharjee, 2015; Ebrahimi & Blackman, 2016; Dhang et al., 2020), can operate with or without the presence of net magnetic flux, and may well be important in magnetized discs. But in discs with a conserved net vertical flux, mean-field effects are not essential to create and maintain persistent toroidal and radial fields: given the net flux and the disc’s mean velocity field (both the shear and radial flow), these fields can be maintained kinematically.
There is also extensive literature adopting the kinematic approach, much of it concerned with connecting the disc’s dynamo field with magnetocentrifugal winds (Königl, 1989; Li, 1995; Ferreira & Pelletier, 1995; Ferreira, 1997). This paper falls in the latter camp, although we focus almost exclusively on the structure of the disc interior and consider the wind primarily as a sink of energy and angular momentum — mainly through its Poynting flux. An unusual feature of our approach is that we abandon the assumption of radial self-similarity, which is adopted in most treatments in the literature. As our objective is to isolate the most robust features of elevated disc dynamos, we create a “minimal" dynamo model that nevertheless admits physically interesting analytic solutions. We ignore all stochastic mean-field effects except for a turbulent resistivity. While we calculate the radial inflow speed self-consistently, we assume Keplerian angular velocity and adopt a simple model for the vertical distribution of density and turbulent pressure.
2.1 Dynamo in a magnetized slab disc
We base our model on the induction equation with a scalar turbulent resistivity $\eta$,
@@ 2.1 Dynamo in a magnetized slab disc
but for simplicity assume that the resistivity only acts on derivatives with respect to $z$ (in cylindrical coordinates). This assumption is generally consistent with a thin disc (scale height $H\ll r$), provided that $|B_{r}|\ga(H/r)B_{z}$ near the disk surface (where we take $B_{z}$ to be positive).
@@
(where partial derivatives with respect to $z$ are denoted by a prime), we only need two components of the simplified induction equation to describe the dynamo, e.g.
@@
The final term on the right-hand side of equation (9) represents the angular momentum loss due to internal viscous stresses within the disc, where ${\rm Pr}_{t}$ is the turbulent magnetic Prandtl number (the ratio of turbulent viscosity to resistivity) and we have adopted the expression for a steady-state Keplerian disc for simplicity. The first term on the right-hand side represents the divergence of the large-scale stress, where we have kept only the term associated with the vertical flux of angular momentum; here $\Omega$ is the Keplerian angular velocity. In the applications below, the term corresponding to the radial flux is never larger than the viscous term and is typically smaller.
@@ 2.2 Boundary conditions
In our simple slab geometry, we assume that the disc has a sharp boundary at $z=H$, where the density and resistivity transition from finite $\rho,\eta$ to zero. This is obviously an artificial assumption made for computational simplicity, but in fact it is physically reasonable. Although there is no wind in our model, the magnetic field exterior to the disc carries both angular momentum and energy away from the disc. Outside the disc, the field should adopt a force-free configuration quantitatively similar to disc-wind models in the limit that the Alfvén point along each streamline lies far outside its footpoint. In the self-similar formulation of Blandford & Payne (1982), this limit corresponds to $\lambda\rightarrow\infty$, $\kappa\rightarrow 0$.
@@ 2.3 Disc evolution equations
Second, the pressure near the midplane must exceed the pinching force exerted by $B_{\phi}$ and $B_{r}$. As emphasized by Ferreira & Pelletier (1993), both of these pressure forces are directed toward the midplane at small $z$, and thus contribute to confining the disc vertically. The scale height of the disc is determined by some internal pressure, $P_{d}$, which could be the usual gas or radiation pressure in the disc interior, or the same turbulent pressure that determines $\eta$. The requirement that gravity ultimately confine the disc, $P_{d}\sim\rho H^{2}\Omega^{2}$, imposes the self-consistency condition that $B_{\phi}^{2}+B_{r}^{2}<8\pi\rho H^{2}\Omega^{2}$ at all $z<H$. A necessary condition, $B_{\phi}^{2}(H)<8\pi\rho H^{2}\Omega^{2}$, gives $\xi^{2}(q\cot q-1)^{2}<8q^{2}/3$. For $\xi\la 1$, this condition permits all values of $20<q<\pi/$ 2 and $\xi$-dependent bands at higher $q$. Additional constraints may restrict the parameter space further.
A steady state is attained for $v_{B}=0$ and $\dot{M}=\rho rH\bar{v}$ independent of radius. Applying the first condition with equation (20), we obtain a mean accretion speed
@@
The magnetic flux is stationary when $\xi q\cot q=(3/2){\rm Pr}_{t}(H/r)$. Since we expect $q\sim O(1)$ under the action of MRI in a standard thin accretion disc with a weak magnetic field (§3), we recover the condition obtained by Lubow et al. (1994) for a thin disc to retain poloidal field, i.e., the field needs to be very weakly inclined with respect to the vertical, $\xi\sim H/r$. However, our model exhibits a novel feature if $q$ approaches $\pi/2$: in this case, $q\cot q\ll 1$ and flux can be retained even if $\xi\sim O(1)$, as is required most large-scale disc-wind models (e.g., Blandford & Payne, 1982; Königl, 1989; Li, 1995; Lubow et al., 1994; Okuzumi et al., 2014; Takeuchi & Okuzumi, 2014). The reason for this is twofold. First, the inward advection speed increases due to large-scale magnetic torques. Second, the vertical gradient of $B_{r}$ decreases near the disc surface, which reduces the outward diffusion. Under these conditions the accretion is driven almost entirely by large-scale magnetic stresses. We refer to such discs as “magnetically boosted."
For $q>\pi/2$, $q\cot q$ becomes negative and we see from equation (20) that resistive diffusion actually helps to trap the flux and draw it inward. This is because the poloidal flux surfaces become convex-outward, and thus exert an inward tension force near the disc surface.
2.4 Resistive and viscous dissipation

### v2

@@ Abstract
We present an analytic, quasi-local model for accretion discs threaded by net, vertical magnetic flux. In a simple slab geometry and ignoring stochastic mean-field dynamo effects, we calculate the large-scale field resulting from the balance between kinematic field amplification and turbulent diffusion. The ability of the disc to accumulate magnetic flux is sensitive to a single parameter dependent on the ratio of the vertical diffusion time to the Alfvén crossing time, and we show how the saturation levels of magnetorotational and other instabilities can govern disc structure and evolution. Under wide-ranging conditions, inflow is governed by large-scale magnetic stresses rather than internal viscous stress. We present models of such “magnetically boosted” discs and show that they lack a radiation pressure-dominated zone. Our model can account for “magnetically elevated” discs as well as instances of midplane outflow and field reversals with height that have been seen in some global simulations. Using the time-dependent features of our model, we find that the incorporation of global transport effects into disc structure can lead to steady or episodic “magnetically arrested discs” (MADs) that maximize the concentration of magnetic flux in their central regions.
@@
pubyear: 2024pagerange: A simple model of globally magnetized accretion discsA.2
@@ 1 Introduction
Magnetic boosting of the inflow speed as well as magnetic elevation can be critical for understanding the dynamics, spectra and stability of accretion discs in a variety of astrophysical systems. Even before these computational trends had been established, the existence of magnetically supported discs had been proposed for phenomenological reasons (Pariev et al., 2003; Blaes et al., 2006; Begelman & Pringle, 2007; Gaburov et al., 2012; Sądowski, 2016; Begelman & Silk, 2017; Dexter & Begelman, 2019). Yet while magnetic elevation seems to be a robust feature of simulated flows, there is to date no clear explanation of how large-scale, organized magnetic field structures are created, and what factors determine their properties.
@@
Even more extreme are so-called magnetically arrested discs (MADs) (Narayan et al., 2003; Igumenshchev, 2008), which have been studied intensively as a type of accretion flow that maximizes the magnetic flux threading a black hole, thus maximizing the potential to extract energy from the black hole’s spin (Tchekhovskoy et al., 2011). While the scaling between the accretion rate and trapped magnetic flux in a MAD was initially derived by equating the ram-pressure of free-fall accretion to the magnetic pressure of the trapped flux (Bisnovatyi-Kogan & Ruzmaikin, 1974, 1976; Narayan et al., 2003) — arguing that the magnetic flux would “arrest" the flow — it has since been recognized that MAD models permit continuous accretion and that the accumulation of magnetic flux in a MAD can extend to much larger radii than the black hole horizon, in effect creating a large-scale magnetosphere (McKinney et al., 2012) co-existent with the accretion flow.
@@
Despite the potential for magnetically elevated and arrested discs to explain a wide range of accretion phenomena, there remains deep uncertainty about how to establish and maintain them in the first place. A large body of work, beginning with van Ballegooijen (1989) and Lubow et al. (1994), suggests that under most conditions the flux distribution will evolve with time. In particular, Lubow et al. (1994) highlighted the challenge faced by thin discs in accumulating and retaining flux, finding that the disc will advect and retain flux only if the poloidal field lines at the disc surface make an angle with the vertical smaller than $\sim Hv_{r}/\eta$, where $H$ is the disc scale height, $v_{r}$ is the inflow speed inside the disc, and $\eta$ is the magnetic diffusivity in the disc interior, in most cases assumed to be due to turbulence. For standard viscous disc models this critical angle is $\sim H/r\ll 1$, whereas constraints imposed by the external electromagnetic environment of the disc — whether by matching the interior field to a magnotocentrifugal wind (Blandford & Payne, 1982; Königl, 1989; Li, 1995) or some other physically motivated field configuration (Lubow et al., 1994; Okuzumi et al., 2014; Takeuchi & Okuzumi, 2014) — tend to favor an angle of order unity.
@@
Various effects may help a disc to retain magnetic flux, including thickening of the disc due to magnetic pressure support, reduced diffusivity in the disc surface layers (Lovelace et al., 2009; Guilet & Ogilvie, 2012, 2013), and enhanced inflow speeds due to large-scale magnetic torques (Blandford & Payne, 1982; Ferreira & Pelletier, 1993; Königl, 1989; Li, 1995). In this paper, we focus on the latter effect, arguing that enhanced inflow speeds may be a natural feature of accretion discs under many conditions, even in the standard thin-disk limit in which the disc is supported by gas pressure and the magnetorotational instability (MRI) operates as usual.
@@
To make this case — and explore its broader implications — we construct and analyze a quasi-local, time-dependent, analytic model for thin discs threaded by a net vertical magnetic field. The plan of the paper is as follows. In §2 we introduce the model (§2.1) and solve for the magnetic field structure using boundary conditions (§2.2) that connect the disc to its external environment. We use these solutions to derive disc evolution equations (§2.3), which show that magnetic stresses are at least as important for driving inflow and magnetic flux evolution as internal viscous stresses, and are often much more important. We refer to these rapidly inflowing discs as “magnetically boosted.” We conclude the section with a discussion of the reduced levels of dissipation expected inside magnetized discs (§2.4), one consequence of which is the absence of any radiation pressure-dominated inner zone (§4.1).
@@
The analysis of §2 reveals the critical importance of a parameter (labeled $q$) which essentially represents the ratio of the diffusion time across the disc scale height to the Alfvén crossing time. Given $B_{z}$, this ratio mainly depends on the level of turbulence inside the disc, driven by magnetorotational and perhaps other instabilities. In §3, we discuss how this ratio might differ in weakly and strongly magnetized discs, while recognizing that the saturation levels of disc turbulence are among the least understood inputs for this or any similar model. Nevertheless, the adoption of a specific closure relation allows us to construct steady-state models of magnetically boosted accretion discs in §4, for both weakly and strongly magnetized limits. We provide a preliminary, local analysis of the stability of these models in an Appendix.
@@
We summarize our results in §6 and expand on two further implications: the possibility of midplane outflow and field reversals within the disc, and the potential relevance of our model to episodic accretion phenomena such as X-ray nova outbursts and changing-look AGN.
@@ 2 Magnetized disc model
Extensive literature exists on accretion disc dynamos, aimed at understanding both the turbulent and organized magnetic fields (e.g., Blackman & Nauman, 2015, and references therein). Turbulent accretion disc dynamos, which have been the subject of numerous numerical and analytic studies (Kato & Yoshizawa, 1995; Vishniac & Brandenburg, 1997; Ogilvie, 2003; Gressel, 2010; Käpylä & Korpi, 2011; Gressel & Pessah, 2015; Squire & Bhattacharjee, 2015; Ebrahimi & Blackman, 2016; Dhang et al., 2020; Mondal & Bhat, 2023), can operate with or without the presence of net magnetic flux, and may well be important in magnetized discs. An alternate, plausible scenario — which we adopt here — is that a large-scale field is imposed on the disc, either at an outer boundary where matter is introduced into the disc or through diffusion from an ambient medium. Sources of net magnetic flux might include the stellar magnetosphere or stellar wind in a mass-transfer binary, or a large-scale field threading the molecular clouds in a galactic nucleus harboring an AGN. If a conserved net vertical flux threads the inner disc, mean-field dynamo growth from turbulence is not essential to the model, which relies solely on advected flux and shear amplification to supply poloidal and toroidal fields. This is not a dynamo in the classic sense that all components of the large-scale field need to be maintained through a self-consistent cycle. Instead, the (primarily vertical) field associated with the net flux is taken to be indestructible over global radial and temporal scales under consideration — it can only be redistributed within the disc radially. Other field components — primarily radial and azimuthal — evolve through the action of shear (the $\Omega$-effect) and inflow inside the disc, while affecting the near-disc environment through vertical diffusion and stress.
There is an extensive literature adopting this kinematic approach, much of it concerned with connecting the disc’s large-scale field with magnetocentrifugal winds (Königl, 1989; Li, 1995; Ferreira & Pelletier, 1995; Ferreira, 1997). This paper falls in this camp, although we focus almost exclusively on the structure of the disc interior and consider the wind primarily as a sink of energy and angular momentum — mainly through its Poynting flux. An unusual feature of our approach is that we abandon the assumption of radial self-similarity, which is adopted in most treatments in the literature. As our objective is to isolate the most robust features of magnetically elevated discs, we create a “minimal" model that nevertheless admits physically interesting analytic solutions. We ignore all stochastic mean-field effects except for a turbulent diffusivity. While we calculate the radial inflow speed self-consistently, we assume Keplerian angular velocity and adopt a simple model for the vertical distribution of density and turbulent pressure.
2.1 Magnetized slab disc
We base our model on the induction equation with a scalar turbulent diffusivity $\eta$,
@@ 2.1 Magnetized slab disc
but for simplicity assume that the diffusivity only acts on derivatives with respect to $z$ (in cylindrical coordinates). This assumption is generally consistent with a thin disc (scale height $H\ll r$), provided that $|B_{r}|\ga(H/r)B_{z}$ near the disk surface (where we take $B_{z}$ to be positive). At this point the source and level of turbulence responsible for $\eta$ are unspecified; in §3 we will discuss how it might be set quantitatively by the magnetorotational instability (MRI).
@@
(where partial derivatives with respect to $z$ are denoted by a prime), we only need two components of the simplified induction equation, e.g.
@@
The final term on the right-hand side of equation (9) represents the angular momentum loss due to internal viscous stresses within the disc, where ${\rm Pr}_{t}$ is the turbulent magnetic Prandtl number (the ratio of turbulent viscosity to magnetic diffusivity) and we have adopted the expression for a steady-state Keplerian disc for simplicity. The first term on the right-hand side represents the divergence of the large-scale stress, where we have kept only the term associated with the vertical flux of angular momentum; here $\Omega$ is the Keplerian angular velocity. In the applications below, the term corresponding to the radial flux is never larger than the viscous term and is typically smaller.
@@ 2.2 Boundary conditions
In our simple slab geometry, we assume that the disc has a sharp boundary at $z=H$, where the density and diffusivity transition from finite $\rho,\eta$ to zero. This is obviously an artificial assumption made for computational simplicity, but in fact it is physically reasonable. Although there is no wind in our model, the magnetic field exterior to the disc carries both angular momentum and energy away from the disc. Outside the disc, the field should adopt a force-free configuration quantitatively similar to disc-wind models in the limit that the Alfvén point along each streamline lies far outside its footpoint. In the self-similar formulation of Blandford & Payne (1982), this limit corresponds to $\lambda\rightarrow\infty$, $\kappa\rightarrow 0$.
@@ 2.3 Disc evolution equations
Second, the pressure near the midplane must exceed the pinching force exerted by $B_{\phi}$ and $B_{r}$. As emphasized by Ferreira & Pelletier (1993), both of these pressure forces are directed toward the midplane at small $z$, and thus contribute to confining the disc vertically. The scale height of the disc is determined by some internal pressure, $P_{d}$, which could be the usual gas or radiation pressure in the disc interior or the same turbulent pressure that determines $\eta$. The requirement that gravity ultimately confine the disc, $P_{d}\sim\rho H^{2}\Omega^{2}$, imposes the self-consistency condition that $B_{\phi}^{2}+B_{r}^{2}<8\pi\rho H^{2}\Omega^{2}$ at all $z<H$. A necessary condition, $B_{\phi}^{2}(H)<8\pi\rho H^{2}\Omega^{2}$, gives $\xi^{2}(q\cot q-1)^{2}<8q^{2}/3$. For $\xi\la 1$, this condition permits all values of $20<q<\pi/$ 2 and $\xi$-dependent bands at higher $q$. Additional constraints may restrict the parameter space further.
A steady state is attained for $v_{B}=0$ and an accretion rate $\dot{M}\propto\rho rH\bar{v}$ independent of radius. Applying the first condition with equation (20), we obtain a mean accretion speed
@@
The magnetic flux is stationary when $\xi q\cot q=(3/2){\rm Pr}_{t}(H/r)$. Since we expect $q\sim O(1)$ under the action of MRI in a standard thin accretion disc with a weak magnetic field (§3), we recover the condition obtained by Lubow et al. (1994) for a thin disc to retain poloidal field, i.e., the field needs to be very weakly inclined with respect to the vertical, $\xi\sim H/r$. However, our model exhibits a novel feature if $q$ approaches $\pi/2$: in this case, $q\cot q\ll 1$ and flux can be retained even if $\xi\sim O(1)$, as is required in most large-scale disc-wind models (e.g., Blandford & Payne, 1982; Königl, 1989; Li, 1995; Lubow et al., 1994; Okuzumi et al., 2014; Takeuchi & Okuzumi, 2014). The reason for this is twofold. First, the inward advection speed increases due to large-scale magnetic torques. Second, the vertical gradient of $B_{r}$ decreases near the disc surface, which reduces the outward diffusion. Under these conditions the accretion is driven almost entirely by large-scale magnetic stresses. We refer to such discs as “magnetically boosted."
For $q>\pi/2$, $q\cot q$ becomes negative and we see from equation (20) that magnetic diffusion actually helps to trap the flux and draw it inward. This is because the poloidal flux surfaces become convex-outward, and thus exert an inward tension force near the disc surface.
2.4 Dissipation